ball with a diameter of 20,00 cm is placed in water with 75% of its volume above the surface of the water. The buoyant force of the water on the ball is…?


1
Expert's answer
2022-06-02T10:25:53-0400
"V=\frac{4}{3}\pi R^3=\frac{4}{3}*3.14* 0.1^3=4.19*10^{-3}\:\rm m^3"

"F_b=g\rho_wV_s=g\rho_w*0.25V\\
=9.8*1000*0.25*4.19*10^{-3}=10\:\rm N"